OpenStudy (anonymous):

Sue initially has 5 hours of pop music and 4 hours of classical music in her collection. Every month onwards, the hours of pop music in her collection is 25% more than what she had the previous month. Her classical music does not change. Which function shows the total hours of music she will have in her collection after x months? f(x) = 4(1.25)x + 5 f(x) = 5(1.25)x + 4 f(x) = 4(0.25)x + 5 f(x) = 5(0.25)x + 4

OpenStudy (anonymous):

@Michele_Laino , @Preetha , @welshfella

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

x is in the exponent in all of the options. Please use a caret (the ^ symbol) to indicate an exponent.

OpenStudy (solomonzelman):

She has 4 hours in collection, and that increases by 25%,. This is: \(\large\color{black}{ \displaystyle y=4(1.25)^x }\) + initial 5 hours: \(\large\color{black}{ \displaystyle y=4(1.25)^x +5 }\)

OpenStudy (anonymous):

Oh so it is the first one A ? @SolomonZelman

OpenStudy (welshfella):

it is the pop music collections which grows each month so it cant be A

OpenStudy (welshfella):

- she starts with 5 hours of pop not 4

OpenStudy (solomonzelman):

Thanks for correcting me.... I read 5 is classic 4 is pop for some reason, dky:)

OpenStudy (solomonzelman):

yes, it should be changed (switch 4 & 5)